l. Assume that j : R-→ R-s C and satisfies what are known as the Cauchy-Riemann equations:
l. Assume that j : R-→ R-s C and satisfies what are known as the Cauchy-Riemann equations:
(c) Let r-(r1, 2) and (s1, s2) be vectors in IR2 and suppose that (ri, 2)f(s1, 82) and Df(81,82)メ0. Show that f-1 satisfies the Cauchy-Riemann equations when evaluated at r. (Hint: Might I make a notational suggestion: Leta(s) = sim) = % (n, s) and b(s) 쓺(81, 82) =-警( )) 81,82 (d) For this last bit, drop the assumption that f satisfies the Cauchy-Riemann equations. Give a function whose derivative is Do for all z E R2. Explain why this f is not invertible (although Df(x)メ0).
Answer #1

a) We are using the suggested notation.

Since Df(s_1,s_2)\neq 0 we know that either a(s)\neq 0 or b(s)\neq 0, and thus, in any case, Df(s_1,s_2)\neq 0 implies

\det Df(s_1,s_2) =a(s)^2+b(s)^2> 0

Since (f^{-1}\circ f )(x,y)=(x,y) for all (x,y)\in\mathbb R^2 , we have

D(f^{-1}\circ f)(x,y)= \begin{pmatrix}1&0\\0&1\end{pmatrix}

for all (x,y)\in\mathbb R^2 . Using chain rule, we get

\begin{align*}Df^{-1}(r)\cdot Df(s)&=D(f^{-1}\circ f)(s)\\ &= \begin{pmatrix}1&0\\0&1\end{pmatrix}\\ \Rightarrow\hspace{2cm}Df^{-1}(r)&=(Df(s))^{-1} \\ &={\frac 1{a(s)^2+b(s)^2}}\begin{pmatrix}a(s)&b(s)\\ -b(s)&a(s)\end{pmatrix} \end{align*}

Since

\begin{align*}Df^{-1}(r)=\begin{pmatrix}{\frac{\partial f^{-1}_1}{\partial x}}(r)&{\frac{\partial f^{-1}_1}{\partial y}}(r)\\ {\frac{\partial f^{-1}_2}{\partial x}}(r)&{\frac{\partial f^{-1}_2}{\partial y}}(r)\end{pmatrix} \end{align*}

we conclude that

\begin{align*}{\frac{\partial f^{-1}_1}{\partial x}}(r)&={\frac{a(s)}{a(s)^2+b(s)^2}}={\frac{\partial f^{-1}_2}{\partial y}}(r)\\ {\frac{\partial f^{-1}_1}{\partial y}}(r)&={\frac{a(s)}{a(s)^2+b(s)^2}}= -{\frac{\partial f^{-1}_2}{\partial x}}(r)\end{align*}

Therefore, \begin{align*}f^{-1}\end{align*} satisfies Cauchy-Riemann.

b) The function \begin{align*}f(x,y)=(x,0)\end{align*} has derivative

\begin{align*}Df(r_1,r_2)&=\begin{pmatrix}{\frac{\partial f_1}{\partial x}}(r)&{\frac{\partial f_1}{\partial y}}(r)\\ {\frac{\partial f_2}{\partial x}}(r)&{\frac{\partial f_2}{\partial y}}(r)\end{pmatrix}\\ &=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}\end{align*}

for all x=(r_1,r_2)\in\mathbb R^2 .

Since f(x,0)=(x,0)=f(x,1) for all x\in\mathbb R, the function \begin{align*}f(x,y)=(x,0)\end{align*} ​​​​​​​ is not injective, hence, not invertible either.